An analyst is using a sample of 15 P/Es to estimate the average P/E for the S&P 500 Index. The sample mean is 16.0 and its variance is 20.0. If the population mean and standard deviation are unknown, then the standard error of the sample mean is closest to:

Group of answer choices

A. 4.47

B. 1.15

C. 0.30
Answer #1

ANSWER:

Standard error = standard deviation / (n) ^ 1/2

std = (variance) ^ 1 / 2 = (20) ^ 1 / 2 = 4.4721

n = 15

standard error = 4.4721 / (15) ^ 1 / 2 = 4.4721 / 3.8729 = 1.1547

so option b is the right answer.